# Katsura’s self-similar groupoid actions, Putnam’s binary factors, and their limit spaces

### v1

@@ Abstract.
We show that the dynamical system associated by Putnam to a pair of graph embeddings is identical to the shift map on the limit space of a self-similar groupoid action on a graph. Moreover, performing a certain out-split on said graph gives rise to a Katsura groupoid action on the out-split graph whose associated limit space dynamical system is conjugate to the previous one. We characterise the self-similar properties of these groupoids in terms of properties of their defining data, two matrices $A$, $B$. We prove a large class of the associated limit spaces are bundles of circles and points which fibre over a totally disconnected space, and the dynamics restricted to each circle is of the form $z\to z^{n}$. Moreover, we find a planar embedding of these spaces, thereby answering a question Putnam posed in his paper.
@@ 1. Introduction
One of the most beautiful aspects of self-similar group theory is its connections, discovered by Nekrashevych [11], to the theory of dynamical systems. To any contracting self-similar group, one can construct its limit dynamical system, which is a self map of a compact metric space whose dynamical properties are governed by the properties of the self-similar group, and vice-versa. Many natural dynamical systems arise as examples, for instance hyperbolic post-critically finite rational maps acting on their Julia sets. This description was used, for instance, to solve the Twisted Rabbit Problem [1].
@@
An embedding pair consists of two directed graphs $H$ and $E$, along with a pair of embeddings $\xi^{0},\xi^{1}:H\hookrightarrow E$ satisfying certain conditions. Putnam’s factor is obtained by identifying two (one-sided) infinite paths in $E$ that arise from the same path in $H$ embedded along two binary sequences of embeddings that are related through carry over in binary addition. He then proves that the natural extension of these dynamical systems are Smale spaces, computes their homology, as well as the $K$-theory of the associated $C^{*}$-algebras. A corollary of his is that these $C^{*}$-algebras exhaust all possible Ruelle algebras arising from irreducible Smale spaces [13, Theorem 6.5].
@@
Using Kitchens’ out-split construction for graphs [9], which we extend to self-similar groupoid actions, we prove that every self-similar groupoid action coming from an embedding pair can be out-split to a Katsura groupoid action. Therefore, Putnam’s dynamical systems are topologically conjugate to the limit dynamical system of a Katsura groupoid action. This explains the similarity in the $K$-theory results [13, Theorem 6.1 and 6.2] and Theorem 3.1 due to Katsura and Exel-Pardo.
@@
Given a self-similar groupoid action on a graph, we show the connected component space of the limit space can be identified with the quotient of the infinite path space of the graph by a natural equivalence relation. We use this description to prove, for a large class of Katsura groupoids, the connected components in their limit spaces are circles and points, analogous to Putnam’s result [13, Corollary 7.7]. We identify the dynamics on the circle components as $z\to z^{n}$, where $n\in\mathbb{N}$ can vary, dependent on the value of $z$ under a natural factor map to a subshift of finite type. Thus, “Katsura systems” exhibit interesting interplay of 00 and $1$ dimensional dynamics.
This description provided the impetus to look for a planar embedding of the limit space for such Katsura groupoid actions, to better understand how the circles and points are configured. The embedding is reminiscent of a solar system trajectory, with planets orbiting a star and moons orbiting the planet, but ad infinitum. As a corollary, we prove that Putnam’s dynamical systems embed into the plane, answering Putnam’s question (see [13, Question 7.10]).
The paper is organised as follows. In Section 2 we provide background on self-similar groupoid actions and their limit dynamical systems. Section 3 introduces Exel and Pardo’s construction of Katsura groupoid actions on graphs, and contains our matrix characterisation of when they are contracting and regular. Section 4 introduces Putnam’s binary factors of subshifts of finite type, and we prove that they are limit dynamical systems of certain self-similar groupoid actions on graphs. Section 5 defines out-splits and uses them, along with the previous result, to show that Putnam’s dynamical systems are topologically conjugate to the limit space dynamical systems of Katsura groupoid actions over certain out-split graphs. Section 6 contains our results on the connected components of limit spaces. The final section proves that regular Katsura systems with $B\in M_{N}(\{0,1\})$ embed into the plane, and hence, so do Putnam’s dynamical systems.
2. Self-similar groupoid actions on graphs and unital Katsura algebras
In this section we describe self-similar groupoid actions on finite directed graphs and their properties. These generalise the notion of a self-similar group introduced by Bartholdi, Grigorchuk, Nekrashevych and others. The main examples we are interested in are the Katsura groupoid actions [8]. Katsura developed a family of Cuntz-Pimsner algebras using two matrices as models for Kirchberg algebras. Exel and Pardo [4, Section 18] realised these as self-similar groupoids acting on a graph [10, Example 7.7].
@@ 2.2. Self-similar actions of groupoids on graphs
Katsura groupoids are a type of self-similar groupoid action on a graph, so we take a few paragraphs to introduce them. For further details see [10] .
@@
If the action of $G$ on $E^{*}$ is range preserving, then $G_{v}\coloneqq\{g\in G\mid d(g)=v\}$ is a group and hence $G$ is a group bundle. In that case we say $(G,E)$ is a self-similar group bundle on $E$. As we will see later, this holds for all Katsura groupoids by definition.
@@
If we replace $G$ in the definition above with a finite set $A$ with $E^{0}\subseteq A$ and retracts $c,d:A\to E^{0}$, then this is the notion of an automaton defined in [10], and one checks that such a pair extends to an action-restriction pair of the free groupoid associated to $(A,c,d)$. Katsura groupoids are, in general, not generated from an automaton but an action-restriction pair defined above.
@@
2.3. Limit spaces of self-similar groupoid actions on graphs
@@
It is shown in [3, Theorem 4.3] that if $(G,E)$ is a contracting and regular self-similar groupoid such that $E$ has no sources, then $(\tilde{\sigma},\mathcal{J}_{G,E})$ is an open, surjective, and metrically expanding local homeomorphism.
@@
3. Katsura groupoid actions on directed graphs
We now consider the example class of action-restriction pairs that is central to this paper, Katsura groupoid actions on a graph. We completely characterise when these are contracting and regular, which turns out to be rather subtle.
Let $N\in\mathbb{N}$ and $A=(A_{ij})$ be a matrix in $M_{N}(\mathbb{N})$ and $B=(B_{ij})$ such that $A_{ij}=0$ implies $B_{ij}=0$. Then $A$ is the adjacency matrix of the graph $E_{A}$ with
@@ 3. Katsura groupoid actions on directed graphs
Exel and Pardo [4] define a group action $\sigma:\mathbb{Z}\times E_{A}\to E_{A}$ and a one-cocycle $\varphi:\mathbb{Z}\times E_{A}\to\mathbb{Z}$ as follows: write $g\in\mathbb{Z}$ multiplicatively as $g=a^{k}$, $k\in\mathbb{Z}$. Then, $\sigma(a^{k},e_{i,j,m})=e_{i,j,\hat{m}}$ and $\varphi(a^{k},e_{i,j,m})=a^{\hat{k}}$, where
@@
(3.2) $(a_{i}^{k},e_{i,j,m})\to(e_{i,j,\hat{m}},a^{\hat{k}}_{j}).$
See [10, Appendix A] for a more careful treatment of these actions and [10, Example 7.7] for a description of the faithful quotient.
@@
Let $N\in\mathbb{N}$ and let $A=(A_{ij})$ be a matrix in $M_{N}(\mathbb{N})$, and $B=(B_{ij})$ a matrix in $M_{N}(\mathbb{Z})$ such that $A$ has no zero rows and $A_{ij}=0$ implies $B_{ij}=0$. Then, the Cuntz-Pimsner algebra $\mathcal{O}(G_{B},E_{A})$ is separable, nuclear, and in the UCT class. The $K$-theory groups of $\mathcal{O}(G_{B},E_{A})$ are given by
@@
Theorem 3.1 outlines several restrictions that can be put on a Katsura groupoid action whose associated Cuntz-Pimsner algebra is a unital Kirchberg algebra. We now consider restrictions that arise on the self-similar groupoid side. It will be helpful to first understand the kernel of a Katsura action $\phi_{A,B}:\mathbb{Z}\times E^{0}_{A}\to\operatorname{PIso}(E^{*}_{A})$.
@@
Denote the corresponding faithful self-similar groupoid action by $(G_{B},E_{A})$. Note that for $\mu\in E^{n}$ such that $B_{\mu}\neq 0$, $k\frac{B_{\mu}}{A_{\mu}}\in\mathbb{Z}$ if and only if $\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}$ divides $k$. So, by Proposition 3.2, the group $(G_{B})_{i}$ is finite if and only if $\text{max}_{\mu\in iE^{*}:B_{\mu}\neq 0}\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}<\infty$ and its cyclic order $o_{i}$ is the smallest $k\in\mathbb{Z}$ such that $\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}$ divides $k$, for all $B_{\mu}\neq 0$. Thus, $o_{i}=\text{lcm}(\{\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}\mid\mu\in iE_{A}^{*}:B_{\mu}\neq 0\})$. We have the following corollary.
@@
Suppose $(G_{B},E_{A})$ is a Katsura groupoid, and define
@@
Definition 3.5.
@@
Let $(A,B)$ be a Katsura pair. Then, $(G_{B},E_{A})$ is contracting if and only if $(G_{B,\infty},E_{A_{\infty}})$ is contracting.
@@
Now, we determine a necessary and sufficient condition for the infinite part of a Katsura pair to be contracting. Suppose $G$ is a finitely generated groupoid with generating set $S=S^{-1}$ and suppose $(G,E)$ is an action-restriction pair. Following Nekrashevych [11, Definition 2.11.9], the contraction coefficient of $(G,E)$ is the quantity:
(3.5) $\rho=\limsup_{n\to\infty}\left(\limsup_{g\in G,\ell(g)\to\infty}\max_{\mu\in d(g)E^{n}}\frac{\ell(g|_{\mu})}{\ell(g)}\right)^{1/n}$
@@
(3.6) $\rho=\limsup_{n\to\infty}\left(\limsup_{m\to\infty}\max_{\mu\in E_{A}^{n}}\frac{\ell(a^{m}_{r(\mu)}|_{\mu})}{m}\right)^{1/n}$
@@
$\rho$ $=\limsup_{n\to\infty}\left(\limsup_{m\to\infty}\max_{\mu\in E_{A}^{n}}\frac{\ell(a^{m}_{r(\mu)}|_{\mu})}{m}\right)^{1/n}$
@@
Let $(A,B)$ be a Katsura pair. Then, the associated self-similar groupoid $(G_{B},E_{A})$ is contracting if and only if $\limsup_{n\to\infty}\left(\max_{\mu\in E_{A,\infty}^{n}}\frac{|B_{\mu}|}{A_{\mu}}\right)^{1/n}<1$.
Definition 3.10.
@@
Proof.
Proposition 3.12.
Let $(A,B)$ be a Katsura pair. Then, $(G_{B},E_{A})$ is regular if and only if $(G_{B,\infty},E_{A_{\infty}})$ and $(G_{B_{<\infty}},E_{A_{<\infty}})$ are regular
@@
Let $(A,B)$ be a Katsura pair such that $(G_{B},E_{A})$ is contracting. Then, $(G_{B},E_{A})$ is regular if and only if $(G_{B_{<\infty}},E_{A_{<\infty}})$ is regular.
@@
We now determine a necessary and sufficient condition for when the finite part of a Katsura pair is regular.
@@
Let $(A,B)$ be a Katsura pair such that $\text{sup}_{\mu\in E^{*}:B_{\mu}\neq 0}\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}<\infty$. Then, $(G_{B},E_{A})$ is regular if and only if there is $K\in\mathbb{N}$ such that for all $\mu\in E^{K}_{A}$ and $\omega\in s(\mu)E^{*}_{A}$, $A_{\omega}$ divides $\frac{B_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}B_{\omega}$.
@@
Let $(A,B)$ be a Katsura pair. Then, the associated self-similar groupoid $(G_{B},E_{A})$ is contracting and regular if and only if $\limsup_{n\to\infty}\left(\max_{\mu\in E_{A,\infty}^{n}}\frac{|B_{\mu}|}{A_{\mu}}\right)^{1/n}<1$ and there is $K\in\mathbb{N}$ such that $A_{\omega}$ divides $\frac{B_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}B_{\omega}$ for all $\mu\in E^{K}_{A,<\infty}$ and $\omega\in s(\mu)E^{*}_{A,<\infty}$
@@
Let $(A,B)$ be a Katsura pair such that $B\in M_{N}(\{0,1\})$. If $(G_{B},E_{A})$ is regular, then $(G_{B},E_{A})$ is contracting. Moreover, $(G_{B},E_{A})$ is regular if and only if there is $K\in\mathbb{N}$ such that
@@
To prove that the two bullet points imply $(G_{B},E_{A})$ is regular, it therefore suffices to show (by Proposition 3.12 and 3.13) that $(G_{B,\infty},E_{A,\infty})$ is contracting, but this is the same argument as above. ∎
@@
Figure 1. The graph $E_{A}$ specified by the adjacency matrix $A$ from Examples 3.18 and 3.19.
@@
Let $(G_{B},E_{A})$ be the Katsura groupoid action defined by
@@
Let $(G_{B},E_{A})$ be the Katsura groupoid action defined by
@@
Let $(G_{B},E_{A})$ be the Katsura groupoid action defined by
@@
Let $(G_{B},E_{A})$ be the Katsura groupoid action defined by

### v2

# Katsura-Exel-Pardo self-similar actions, Putnam’s binary factors, and their limit spaces

@@ Abstract.
We show that the dynamical system associated by Putnam to a pair of graph embeddings is identical to the shift map on the limit space of a self-similar groupoid action on a graph. Moreover, performing a certain out-split on said graph gives rise to a Katsura-Exel-Pardo groupoid action on the out-split graph whose associated limit space dynamical system is conjugate to the previous one. We characterise the self-similar properties of these groupoids in terms of properties of their defining data, two matrices $A$, $B$. We prove a large class of the associated limit spaces are bundles of circles and points which fibre over a totally disconnected space, and the dynamics restricted to each circle is of the form $z\to z^{n}$. Moreover, we find a planar embedding of these spaces, thereby answering a question Putnam posed in his paper.
@@
We’d like to thank Enrique Pardo for comments on an early draft.
@@ 1. Introduction
One of the most beautiful aspects of self-similar group theory are its connections, discovered by Nekrashevych [11], to the theory of dynamical systems. To any contracting self-similar group, one can construct its limit dynamical system, which is a self map of a compact metric space whose dynamical properties are governed by the properties of the self-similar group, and vice-versa. Many natural dynamical systems arise as examples, for instance hyperbolic post-critically finite rational maps acting on their Julia sets. This description was used, for instance, to solve the Twisted Rabbit Problem [1].
@@
An embedding pair consists of two directed graphs $H$ and $E$, along with a pair of embeddings $\xi^{0},\xi^{1}:H\hookrightarrow E$ satisfying certain conditions. Putnam’s factor is obtained by identifying two (one-sided) infinite paths in $E$ that arise from the same path in $H$ embedded along two binary sequences of embeddings that are related through carry over in binary addition. He then proves that the natural extension of these dynamical systems are Smale spaces, computes their homology, as well as the $K$-theory of the associated $C^{*}$-algebras. As a corollary, Putnam proves that these $C^{*}$-algebras exhaust all possible Ruelle algebras arising from irreducible Smale spaces [13, Theorem 6.5].
@@
Using Kitchens’ out-split construction for graphs [9], which we extend to self-similar groupoid actions, we prove that every self-similar groupoid action coming from an embedding pair can be out-split to a Katsura-Exel-Pardo action. Therefore, Putnam’s dynamical systems are topologically conjugate to the limit dynamical system of a Katsura-Exel-Pardo action. This explains the similarity in the $K$-theory results [13, Theorem 6.1 and 6.2] and Theorem 3.1 due to Katsura and Exel-Pardo.
@@
Given a self-similar groupoid action on a graph, we show the connected component space of the limit space can be identified with the quotient of the infinite path space of the graph by a natural equivalence relation. We use this description to prove, for a large class of Katsura-Exel-Pardo actions, the connected components in their limit spaces are circles and points, analogous to Putnam’s result [13, Corollary 7.7]. We identify the dynamics on the circle components as $z\to z^{n}$, where $n\in\mathbb{N}$ can vary, dependent on the value of $z$ under a natural factor map to a subshift of finite type. Thus, “Katsura-Exel-Pardo systems” exhibit interesting interplay of 00 and $1$ dimensional dynamics.
This description provided the impetus to look for a planar embedding of the limit space for such Katsura-Exel-Pardo actions, to better understand how the circles and points are configured. The embedding is reminiscent of a solar system trajectory, with planets orbiting a star and moons orbiting the planet, but ad infinitum. As a corollary, we prove that Putnam’s dynamical systems embed into the plane, answering Putnam’s question (see [13, Question 7.10]).
The paper is organised as follows. In Section 2 we provide background on self-similar groupoid actions and their limit dynamical systems. Section 3 introduces Katsura’s construction and Exel and Pardo’s realisation of these as Katsura-Exel-Pardo groupoid actions on graphs, and contains our matrix characterisation of when they are contracting and regular. Section 4 introduces Putnam’s binary factors of subshifts of finite type, and we prove that they are limit dynamical systems of certain self-similar groupoid actions on graphs. Section 5 defines out-splits and uses them, along with the previous result, to show that Putnam’s dynamical systems are topologically conjugate to the limit space dynamical systems of Katsura-Exel-Pardo actions over certain out-split graphs. Section 6 contains our results on the connected components of limit spaces. The final section proves that regular Katsura-Exel-Pardo systems with $B\in M_{N}(\{0,1\})$ embed into the plane, and hence, so do Putnam’s dynamical systems.
2. Self-similar groupoid actions on graphs
In this section we describe self-similar groupoid actions on finite directed graphs and their properties. These generalise the notion of a self-similar group introduced by Bartholdi, Grigorchuk, Nekrashevych and others.
@@ 2.2. Self-similar actions of groupoids on graphs
Katsura-Exel-Pardo actions are a type of self-similar groupoid action on a graph, so we take a few paragraphs to introduce them. For further details see [10] .
@@
If the action of $G$ on $E^{*}$ is range preserving, then $G_{v}\coloneqq\{g\in G\mid d(g)=v\}$ is a group and hence $G$ is a group bundle. In that case we say $(G,E)$ is a self-similar group bundle on $E$. As we will see later, this holds for all Katsura-Exel-Pardo actions by definition.
@@
If we replace $G$ in the definition above with a finite set $A$ with $E^{0}\subseteq A$ and retracts $c,d:A\to E^{0}$, then this is the notion of an automaton defined in [10], and one checks that such a pair extends to an action-restriction pair of the free groupoid associated to $(A,c,d)$. Katsura-Exel-Pardo actions are, in general, not generated from an automaton but an action-restriction pair defined above.
@@
2.3. Properties and limit spaces of self-similar groupoid actions on graphs
In this section we recall standard properties and constructions associated to action-restriction pairs and self-similar group(oid)s acting on the path space of a graph.
Definition 2.4.
Definition 2.5.
Proposition 2.6.
Proof.
@@ 2.3. Properties and limit spaces of self-similar groupoid actions on graphs
It is shown in [3, Theorem 4.3] that if $(G,E)$ is a contracting and regular self-similar groupoid such that $E$ has no sources, then $(\tilde{\sigma},\mathcal{J}_{G,E})$ is an open, surjective, and positively expansive local homeomorphism.
@@
3. Katsura-Exel-Pardo groupoid actions on directed graphs
The main examples of self-similar groupoid actions on graphs that we are interested in are the Katsura-Exel-Pardo groupoid actions [8]. Katsura developed a family of Cuntz-Pimsner algebras using two matrices as models for Kirchberg algebras. Exel and Pardo [4, Section 18] realised these as self-similar groupoids acting on a graph [10, Example 7.7]. For Katsura-Exel-Pardo actions we completely characterise when these are contracting and regular, which turns out to be rather subtle.
Let $N\in\mathbb{N}$. A Katsura pair is a matrix $A=(A_{ij})$ in $M_{N}(\mathbb{N})$ and a matrix $B=(B_{ij})$ such that $A_{ij}=0$ implies $B_{ij}=0$. Then $A$ is the adjacency matrix of the graph $E_{A}$ with
@@ 3. Katsura-Exel-Pardo groupoid actions on directed graphs
Exel and Pardo [4] realised that a Katsura pair gives a self-similar action on a graph in the following way. Define a group action $\sigma:\mathbb{Z}\times E_{A}\to E_{A}$ and a one-cocycle $\varphi:\mathbb{Z}\times E_{A}\to\mathbb{Z}$ as follows: write $g\in\mathbb{Z}$ multiplicatively as $g=a^{k}$, $k\in\mathbb{Z}$. Then, $\sigma(a^{k},e_{i,j,m})=e_{i,j,\hat{m}}$ and $\varphi(a^{k},e_{i,j,m})=a^{\hat{k}}$, where
@@
(3.2) $(a_{i}^{k},e_{i,j,m})\to(e_{i,j,\hat{m}},a^{\hat{k}}_{j}),$
and we call these Katsura-Exel-Pardo groupoid actions (KEP-actions). See [10, Appendix A] for a more careful treatment of these actions and [10, Example 7.7] for a description of the faithful quotient. We reserve the notation $(G_{B},E_{A})$ for the corresponding faithful KEP-action associated with a Katsura pair of matrices $A$ and $B$.
@@
Let $N\in\mathbb{N}$ and let $A=(A_{ij})$ be a matrix in $M_{N}(\mathbb{N})$, and $B=(B_{ij})$ a matrix in $M_{N}(\mathbb{Z})$ such that $A$ has no zero rows and $A_{ij}=0$ implies $B_{ij}=0$. Then, the Cuntz-Pimsner algebra of the associated self-similar groupoid action $\mathcal{O}(G_{B},E_{A})$ is separable, nuclear, and in the UCT class. The $K$-theory groups of $\mathcal{O}(G_{B},E_{A})$ are given by
@@
Theorem 3.1 outlines several restrictions that can be put on a KEP-action whose associated Cuntz-Pimsner algebra is a unital Kirchberg algebra. We now consider restrictions that arise on the self-similar groupoid side. It will be helpful to first understand the kernel of a KEP-action $\phi_{A,B}:\mathbb{Z}\times E^{0}_{A}\to\operatorname{PIso}(E^{*}_{A})$.
@@
Note that for $\mu\in E^{n}$ such that $B_{\mu}\neq 0$, $k\frac{B_{\mu}}{A_{\mu}}\in\mathbb{Z}$ if and only if $\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}$ divides $k$. So, by Proposition 3.2, the group $(G_{B})_{i}$ is finite if and only if $\text{max}_{\mu\in iE^{*}:B_{\mu}\neq 0}\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}<\infty$ and its cyclic order $o_{i}$ is the smallest $k\in\mathbb{Z}$ such that $\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}$ divides $k$, for all $B_{\mu}\neq 0$. Thus, $o_{i}=\text{lcm}(\{\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}\mid\mu\in iE_{A}^{*}:B_{\mu}\neq 0\})$. We have the following corollary.
@@
Suppose $(G_{B},E_{A})$ is a KEP-action, and define
@@
For the following proposition, recall the notion of contracting from Section 2.3.
@@
Let $(A,B)$ be a Katsura pair. Then, the KEP-action $(G_{B},E_{A})$ is contracting if and only if $(G_{B,\infty},E_{A_{\infty}})$ is contracting.
@@
Now, we determine a necessary and sufficient condition for the infinite part of a KEP-action to be contracting. Suppose $G$ is a finitely generated groupoid with generating set $S=S^{-1}$, with associated length function $\ell_{S}:G\to\mathbb{N}$, and suppose $(G,E)$ is an action-restriction pair. Following Nekrashevych [11, Definition 2.11.9], the contraction coefficient of $(G,E)$ is the quantity:
(3.5) $\rho=\limsup_{n\to\infty}\left(\limsup_{g\in G,\ell_{S}(g)\to\infty}\max_{\mu\in d(g)E^{n}}\frac{\ell_{S}(g|_{\mu})}{\ell_{S}(g)}\right)^{1/n}.$
@@
(3.6) $\rho=\limsup_{n\to\infty}\left(\limsup_{m\to\infty}\max_{\mu\in E_{A}^{n}}\frac{\ell_{S}(a^{m}_{r(\mu)}|_{\mu})}{m}\right)^{1/n}$
@@
$\rho$ $=\limsup_{n\to\infty}\left(\limsup_{m\to\infty}\max_{\mu\in E_{A}^{n}}\frac{\ell_{S}(a^{m}_{r(\mu)}|_{\mu})}{m}\right)^{1/n}$
@@
Let $(A,B)$ be a Katsura pair. Then, the associated KEP-action $(G_{B},E_{A})$ is contracting if and only if $\limsup_{n\to\infty}\left(\max_{\mu\in E_{A,\infty}^{n}}\frac{|B_{\mu}|}{A_{\mu}}\right)^{1/n}<1$.
For the following proposition, recall the notion of regular from Section 2.3.
@@
Let $(A,B)$ be a Katsura pair. Then, the KEP-action $(G_{B},E_{A})$ is regular if and only if $(G_{B,\infty},E_{A_{\infty}})$ and $(G_{B_{<\infty}},E_{A_{<\infty}})$ are regular.
@@
Let $(A,B)$ be a Katsura pair such that the KEP-action $(G_{B},E_{A})$ is contracting. Then, $(G_{B},E_{A})$ is regular if and only if $(G_{B_{<\infty}},E_{A_{<\infty}})$ is regular.
@@
We now determine a necessary and sufficient condition for when the finite part of a KEP-action is regular.
@@
Let $(A,B)$ be a Katsura pair such that $\text{sup}_{\mu\in E^{*}:B_{\mu}\neq 0}\frac{A_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}<\infty$. Then, the KEP-action $(G_{B},E_{A})$ is regular if and only if there is $K\in\mathbb{N}$ such that for all $\mu\in E^{K}_{A}$ and $\omega\in s(\mu)E^{*}_{A}$, $A_{\omega}$ divides $\frac{B_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}B_{\omega}$.
@@
Let $(A,B)$ be a Katsura pair. Then, the KEP-action $(G_{B},E_{A})$ is contracting and regular if and only if $\limsup_{n\to\infty}\left(\max_{\mu\in E_{A,\infty}^{n}}\frac{|B_{\mu}|}{A_{\mu}}\right)^{1/n}<1$ and there is $K\in\mathbb{N}$ such that $A_{\omega}$ divides $\frac{B_{\mu}}{\text{gcd}(A_{\mu},|B_{\mu}|)}B_{\omega}$ for all $\mu\in E^{K}_{A,<\infty}$ and $\omega\in s(\mu)E^{*}_{A,<\infty}$
@@
Let $(A,B)$ be a Katsura pair such that $B\in M_{N}(\{0,1\})$. If the KEP-action $(G_{B},E_{A})$ is regular, then $(G_{B},E_{A})$ is contracting. Moreover, $(G_{B},E_{A})$ is regular if and only if there is $K\in\mathbb{N}$ such that
@@
To prove that the two bullet points imply $(G_{B},E_{A})$ is regular, it therefore suffices to show (by Proposition 3.9 and 3.10) that $(G_{B,\infty},E_{A,\infty})$ is contracting, but this is the same argument as above. ∎
@@
Figure 1. The graph $E_{A}$ specified by the adjacency matrix $A$ from Examples 3.15 and 3.16.
@@
Let $(G_{B},E_{A})$ be the KEP-action defined by
@@
Let $(G_{B},E_{A})$ be the KEP-action defined by
@@
Let $(G_{B},E_{A})$ be the KEP-action defined by
@@
Let $(G_{B},E_{A})$ be the KEP-action defined by